# Flow-based linear embedding for Bayesian filtering of nonlinear stochastic dynamical systems

### v1

@@ Abstract
Bayesian filtering for high-dimensional nonlinear stochastic dynamical systems is a fundamental yet challenging problem in many fields of science and engineering. Existing methods face significant obstacles: Gaussian-based filters struggle with non-Gaussian distributions, sequential Monte Carlo methods are computationally intensive and prone to particle degeneracy in high dimensions, and deep learning approaches often fail to balance accuracy and efficiency in complex filtering tasks. To address these challenges, we propose a flow-based Bayesian filter (FBF) that integrates normalizing flows to construct a latent linear state-space model with Gaussian filtering distributions. This framework enables efficient density estimation and sampling through invertible transformations provided by normalizing flows, which can be learned directly from data, thereby eliminating the need for prior knowledge of system dynamics or observation models. Numerical experiments demonstrate the advantages of FBF in terms of both accuracy and efficiency.
@@
Bayesian filtering , Linear embedding , Generative models , Normalizing flows , High-dimensional stochastic systems , Latent state-space modeling
journal: Elsevier
@@ 1 Introduction
Filtering problem is a crucial research area in data assimilation [1], with applications that span numerical weather prediction, oceanography, signal processing, finance, etc. [2, 3]. The primary objective of filtering is to estimate the current state of a dynamical system by integrating noisy observations with prior model information. However, several significant challenges arise when applying filtering methods to more general and complex scenarios, such as nonlinearity and the curse of dimensionality. Developing an efficient computation model for accurate and stable state estimation is of substantial significance and has attracted great attention from scientists and engineers.
@@
Classical filtering methods encompass variational methods and Bayesian filtering. Variational methods, such as 3DVAR [4] and 4DVAR [5], aim to find a best point-estimate of the system’s state by minimizing the misfit for predicting measurements while penalizing deviations from the prior forecast. Bayesian filtering [6], on the other hand, constructs the posterior distribution of the model state of interest through Bayesian principles, thereby effectively quantifying the uncertainties arising from prior information and observations. However, some well-known Bayesian techniques, such as the Kalman filter (KF) [7] and its extensions, including the extended Kalman filter [8] and the ensemble Kalman filter [9, 10, 11, 12], rely on a Gaussian ansatz, rendering them less suitable for applications involving non-Gaussian filtering distributions [13]. Therefore, many researchers utilized the sequential Monte Carlo method, referred to as the particle filter (PF), to obtain a more accurate estimate of the target posterior distribution [14, 15, 16]. Nevertheless, in cases where the system involves complex nonlinearity, PF may require a large number of particles to represent the state accurately, leading to high computational costs. An alternative to these methods is the quadrature Kalman filter (QKF) and its variants [17, 18, 19, 20], which approximate nonlinear systems using quadrature rules. The QKF can offer more accurate estimates for non-Gaussian distributions while generally being less computationally expensive than PF. However, the computational cost of QKF increases significantly as the state dimension grows, making it less suitable for real-time applications in large-scale systems. To overcome this issue, Bao et al. proposed a novel score-based nonlinear filtering methodology [21, 22], which has achieved superior tracking accuracy on extremely high-dimensional stochastic dynamical systems.
@@
The commonality among these aforementioned filters lies in their dependence on precise knowledge and accurate modeling of the underlying dynamics. In contrast, model-free data assimilation techniques, often referred to as data-driven methods, have garnered attention in scenarios where explicit dynamical models are unavailable. Instead, these techniques leverage simulation or experimental data that capture the system’s dynamics to implement state estimation [23, 24, 25, 26, 27]. For example, recurrent neural network (RNN) and its variants, such as long short-term memory network (LSTM) [28] and gated recurrent unit (GRU) [29] have demonstrated remarkable performance in time series-related tasks. However, they often struggle to capture the uncertainty [30]. To this end, there have been significant advances in using probabilistic surrogate modeling techniques to perform state estimation with uncertainty quantification (UQ) [31, 30, 32]. For example, a Kalman filtering-aided encoder-decoder model, termed the recurrent Kalman network (RKN), was proposed in [30]. Similarly, Yuta et al. [32] introduced a variational autoencoder (VAE)-based Bayesian filter that integrates a latent linear Gaussian state transition model, enabling efficient inference of target filtering distributions. Despite these advancements, the aforementioned approaches rely on approximating the target distribution using tractable distributions (e.g., Gaussian distributions), which limits their applicability to scenarios involving complex filtering distributions.
@@ 5 Conclusion
In this paper, a novel flow-based Bayesian filter is proposed to efficiently solve nonlinear filtering problems. In contrast to existing deep learning techniques, the proposed method offers tractable estimation of the target filtering distribution, enabling efficient sampling and high-accuracy density evaluation. Numerical experiments demonstrate the advantages of our framework in terms of accuracy and efficiency, as well as its suitability for addressing high-dimensional filtering problems.
In future work, several questions remain that require further investigation. Firstly, we have focused on establishing the latent space in a data-driven context; however, in practice, partial physical knowledge and sparse data of the underlying dynamical systems may also be available. This motivates the need to explore a physics-informed FBF method. In addition, we will study the extension of the proposed methods for parameter identification within physical models, with the goal of developing a more general flow-based linear embedding Bayesian inference framework. Finally, we aim to enhance the applicability of our approach to real-world scenarios, such as geophysical inversion and climate modeling.

### v2

# Flow-based Bayesian filtering for high-dimensional nonlinear stochastic dynamical systems

@@ Abstract
Bayesian filtering for high-dimensional nonlinear stochastic dynamical systems is a fundamental yet challenging problem in many fields of science and engineering. Existing methods face significant obstacles: Gaussian-based filters struggle with non-Gaussian distributions, while sequential Monte Carlo methods are computationally intensive and prone to particle degeneracy in high dimensions. Although generative models in machine learning have made significant progress in modeling high-dimensional non-Gaussian distributions, their inefficiency in online updating limits their applicability to filtering problems. To address these challenges, we propose a flow-based Bayesian filter (FBF) that integrates normalizing flows to construct a novel latent linear state-space model with Gaussian filtering distributions. This framework facilitates efficient density estimation and sampling using invertible transformations provided by normalizing flows, and it enables the construction of filters in a data-driven manner, without requiring prior knowledge of system dynamics or observation models. Numerical experiments demonstrate the superior accuracy and efficiency of FBF.
@@
Bayesian filtering , Generative models , Normalizing flows , High-dimensional stochastic systems , Latent state-space modeling
journal: arXiv
@@ 1 Introduction
Filtering problem is a crucial research area in data assimilation [1], with applications that span numerical weather prediction, oceanography, signal processing, finance, etc. [2, 3, 4]. The primary objective of filtering is to estimate the current state of a dynamical system by integrating noisy observations with prior model information. However, several significant challenges arise when applying filtering methods to more general and complex scenarios, such as nonlinearity and the curse of dimensionality. Developing an efficient computation model for accurate and stable state estimation is of substantial significance and has attracted great attention from scientists and engineers.
@@
Classical filtering methods encompass variational methods and Bayesian filtering. Variational methods [5, 6], such as 3D-VAR [7] and 4D-VAR [8], aim to find a best point-estimate of the system’s state by minimizing the misfit for predicting measurements while penalizing deviations from the prior forecast. Bayesian filtering [9], on the other hand, constructs the posterior distribution of the model state of interest through Bayesian principles, thereby effectively quantifying the uncertainties arising from prior information and observations. However, some well-known Bayesian techniques, such as the Kalman filter (KF) [10] and its extensions, including the extended Kalman filter [11] and the ensemble Kalman filter [12, 13, 14, 15], rely on a Gaussian ansatz, rendering them less suitable for applications involving non-Gaussian filtering distributions [16]. Therefore, many researchers utilized the sequential Monte Carlo method, referred to as the particle filter (PF), to obtain a more accurate estimate of the target posterior distribution [17, 18, 19]. Nevertheless, in cases where the system involves complex nonlinearity, PF may require a large number of particles to represent the state accurately, leading to high computational costs. An alternative to these methods is the quadrature Kalman filter (QKF) and its variants [20, 21, 22, 23], which approximate nonlinear systems using quadrature rules. The QKF can offer more accurate estimates for non-Gaussian distributions while generally being less computationally expensive than PF. However, the computational cost of QKF increases significantly as the state dimension grows, making it less suitable for real-time applications in large-scale systems. To overcome this issue, Bao et al. proposed a novel score-based nonlinear filtering methodology [24, 25], which has achieved superior tracking accuracy on extremely high-dimensional stochastic dynamical systems.
@@
The commonality among these aforementioned filters lies in their dependence on precise knowledge and accurate modeling of the underlying dynamics. In contrast, model-free data assimilation techniques, often referred to as data-driven methods, have garnered attention in scenarios where explicit dynamical models are unavailable. Instead, these techniques leverage simulation or experimental data that capture the system’s dynamics to implement state estimation [26, 27, 28, 29, 30]. For example, recurrent neural network (RNN) and its variants, such as long short-term memory network (LSTM) [31, 32] and gated recurrent unit (GRU) [33] have demonstrated remarkable performance in time series-related tasks. However, they often struggle to capture the uncertainty [34]. To this end, there have been significant advances in using probabilistic surrogate modeling techniques to perform state estimation with uncertainty quantification (UQ) [35, 34, 36]. For example, a Kalman filtering-aided encoder-decoder model, termed the recurrent Kalman network (RKN), was proposed in [34]. Similarly, Yuta et al. [36] introduced a variational autoencoder (VAE)-based Bayesian filter that integrates a latent linear Gaussian state transition model, enabling efficient inference of target filtering distributions. Despite these advancements, the aforementioned approaches rely on approximating the target distribution using tractable distributions (e.g., Gaussian distributions), which limits their applicability to scenarios involving complex filtering distributions.
@@ 5 Conclusion
In this paper, a novel flow-based Bayesian filter is proposed to efficiently solve nonlinear filtering problems. In contrast to existing deep learning techniques, the proposed method offers tractable estimation of the target filtering distribution, enabling efficient sampling and high-accuracy density evaluation. Numerical experiments demonstrate the advantages of our framework in terms of accuracy and efficiency, as well as its suitability for addressing high-dimensional filtering problems. In future work, we will incorporate physical information into FBF to broaden its applicability and extend its use to real-world scenarios, including geophysical inversion and climate modeling.